# Adaptive tuning of Hamiltonian Monte Carlo methods

### v2

@@ Abstract
With the recently increased interest in probabilistic models, the efficiency of an underlying sampler becomes a crucial consideration. A Hamiltonian Monte Carlo (HMC) is one popular option for models of this kind. Performance of HMC, however, strongly relies on a choice of parameters associated with an integration method for Hamiltonian equations, which up to date remains mainly heuristic or introduces time complexity. We propose a novel computationally inexpensive and flexible approach (we call it Adaptive Tuning or ATune) that, by combining a theoretical analysis of the multivariate Gaussian model with simulation data generated during a burn-in stage of HMC, detects a system specific splitting integrator with a set of reliable HMC hyperparameters, including their credible randomization intervals, to be readily used in a production simulation. The method automatically eliminates those values of simulation parameters which could cause undesired extreme scenarios, such as resonance artefacts, low accuracy or poor sampling. The new approach is implemented in the in-house software package HaiCS, with no computational overheads introduced in a production simulation, and can be easily incorporated in any package for Bayesian inference with HMC. The tests on popular statistical models reveal the superiority of adaptively tuned HMC and generalized Hamiltonian Monte Carlo (GHMC) in terms of stability, performance and accuracy over conventional HMC tuned heuristically and coupled with the well-established integrators. We also claim that the generalized formulation of HMC, i.e. GHMC, is preferable for achieving high sampling performance. The efficiency of the new methodology is assessed in comparison with state-of-the-art samplers, e.g. the No-U-Turn-Sampler, in real-world applications, such as endocrine therapy resistance in cancer, modeling of cell-cell adhesion dynamics and influenza A epidemic outbreak.
@@ 4.1 Randomization interval
GHMC s-AIA3 $\mathcal{U}\{1,...,1332\}$ $\mathcal{U}(0.036,0.052)$ $\mathcal{U}(0,0.5)$
@@
German HMC s-AIA3 $\mathcal{U}\{1,...,16\}$ $\mathcal{U}(0.111,0.160)$ 1
@@
GHMC s-AIA3 $\mathcal{U}\{1,...,7\}$ $\mathcal{U}(0.111,0.160)$ $\mathcal{U}(0,0.5)$
@@
Musk HMC s-AIA3 $\mathcal{U}\{1,...,110\}$ $\mathcal{U}(0.083,0.11)$ 1
@@
GHMC s-AIA3 $\mathcal{U}\{1,...,55\}$ $\mathcal{U}(0.084,0.121)$ $\mathcal{U}(0,0.5)$
@@ 4.2 Validation
The remaining HMC parameters, namely, a number of integration steps per Monte Carlo iteration, $L$, and numbers of burn-in and production iterations, are kept the same as in the numerical experiments in sAIA_paper. In addition, we conduct GHMC simulations to evaluate the efficacy of the proposed randomization interval with this sampler in comparison with HMC. For the PMU in Eq. (9), we set $\varphi$ to be uniformly selected from the interval $(0,0.5)$ (a standard recommendation) and use twice shorter trajectories than for HMC. We remark that this choice of GHMC parameters is based on heuristic intuition. The optimal selection of these parameters are investigated in the following sections.
@@ 8.2 Benchmark tests
Banana $\mathcal{U}\{2,5,7\}$ $\mathcal{U}(0.21904,1)$
@@
G1000 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 39373 2387 1491 2509 5021
meanESS 0.4596 532.1 1071 2142 4285
multiESS 0.3787 299.8 402.5 895.4 3069
1 (HMC) minESS 132650 46712 21620 11740 5554
meanESS 2.530 533.6 1054 2093 4198
multiESS 3.712 301.3 1001 2000 4001
G500 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 31792 1553 919.8 1090 2339
meanESS 0.3429 260.0 520.7 1043 2093
multiESS 0.2553 170.6 320.8 384.0 1422
1 (HMC) minESS 116471 42111 21821 10094 4790
meanESS 2.363 261.3 515.5 1035 2046
multiESS 3.666 107.8 321.6 563.3 1325
G2000 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 63451 3305 2515 5036 9195
meanESS 0.6518 1104 2207 4417 8853
multiESS 0.5029 428.5 860.1 1723 7096
1 (HMC) minESS 174435 47859 22770 10870 9582
meanESS 2.791 1099 2178 4348 8671
multiESS 3.739 677.3 2000 4001 8000
German $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 1.734 13.94 25.93 53.38 113.6
meanESS 0.3463 10.15 23.84 49.05 101.0
multiESS 0.1218 10.62 25.10 49.47 100.5
1 (HMC) minESS 24.92 13.25 26.18 53.09 112.6
meanESS 5.860 10.56 24.45 49.23 102.9
multiESS 2.334 10.88 25.44 49.44 100.8
Musk $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 216.8 143.7 229.4 389.7 601.3 1458
meanESS 145.4 52.08 68.44 165.0 464.5 1031
multiESS 19.79 12.79 36.00 160.4 342.0 774.5
1 (HMC) minESS 6026 1156 187.7 290.3 525.7 1215
meanESS 3068 705.6 143.5 176.3 454.6 1047
multiESS 375.8 72.99 50.60 159.5 279.5 506.9
@@
AT-GHMC $\mathcal{U}(0.152,0.219)$ $\mathcal{U}(2,8)$ $\mathcal{U}(0.21904,1)$ 0.92
HMC tuned $0.111$ 77 1 0.69
HMC $0.05$ 22 11 0.91
$0.5$ 0.02
@@ 9.1 Patient resistance to endocrine therapy in breast cancer
NUTS 100 $\overline{L}=14.8$ $\overline{\Delta t}=0.247$ -
@@
NUTS 100 $\overline{L}=11.8$ $\overline{\Delta t}=0.354$ -
@@
NUTS 100 $\overline{L}=8.1$ $\overline{\Delta t}=0.474$ -
@@ 9.3 Influenza A (H1N1) epidemics outbreak
As an illustrative example, we consider an outbreak of influenza A (H1N1) in 1978 at a British boarding school involving 763 students. The dataset, available in the R package outbreaks outbreaks20, records the daily number of students confined to bed over a 14-day period. Following Grinsztajn2020, we model the outbreak dynamics using a classical SIR (Susceptible-Infected-Recovered) model SIR_paper, coupled with a negative binomial observation model to account for overdispersion in the reported case counts. The model includes the transmission rate, $\beta$, and the recovery rate, $\gamma$, which govern the transitions from susceptible to infected and from infected to recovered, respectively. For the prior distributions, we assume $\beta\sim\mathcal{N}^{0}(1,1)$ and $\gamma\sim\mathcal{N}^{0}(0.4,0.5)$, where $\mathcal{N}^{a}$ denotes a normal distribution truncated at $a$. The prior specification is completed by assigning an exponential prior to the inverse of the dispersion parameter, $\phi$, of the negative binomial distribution, i.e. $1/\phi\sim\text{Exp}(5)$. The initial conditions are set as $I_{0}=1$, $R_{0}=0$, and thus $S_{0}=762$.
@@
NUTS 140 $\overline{L}=5.45$ $\overline{\Delta t}=0.607$ -
@@ Appendix D Sensitivity analysis: Optimal $h_{\text{lower}}$
G1000 HMC (15) 5554 4198 4001
(15)  $5\%$ 5034 4361 4001
(15) + $5\%$ 5107 4483 4001
GHMC (15) 2518 2081 2000
(15)  $5\%$ 2818 2189 2000
(15) + $5\%$ 2772 2231 2000
German HMC (15) 26.18 24.45 25.44
(15)  $5\%$ 25.25 24.77 25.23
(15) + $5\%$ 25.39 25.00 25.30
GHMC (15) 12.38 10.05 10.65
(15)  $5\%$ 11.39 10.18 10.56
(15) + $5\%$ 11.25 10.18 10.65

### v3

@@ Abstract
With the recently increased interest in probabilistic models, the efficiency of an underlying sampler becomes a crucial consideration. Hamiltonian Monte Carlo is one popular option for models of this kind. Performance of the method, however, strongly relies on a choice of parameters associated with an integration for Hamiltonian equations. Up to date, such a choice remains mainly heuristic or introduces time complexity. We propose a novel computationally inexpensive and flexible approach (we call it Adaptive Tuning or ATune) that, by combining a theoretical analysis of the multivariate Gaussian model with simulation data generated during a burn-in stage of a Hamiltonian Monte Carlo simulation, detects a system specific splitting integrator with a set of reliable sampler’s hyperparameters, including their credible randomization intervals, to be readily used in a production simulation. The method automatically eliminates those values of simulation parameters which could cause undesired extreme scenarios, such as resonance artifacts, low accuracy or poor sampling. The new approach is implemented in the in-house software package HaiCS, with no computational overheads introduced in a production simulation, and can be easily incorporated in any package for Bayesian inference with Hamiltonian Monte Carlo. The tests on popular statistical models reveal the superiority of adaptively tuned standard and generalized Hamiltonian Monte Carlo methods in terms of stability, performance and accuracy over conventional Hamiltonian Monte Carlo tuned heuristically and coupled with the well-established integrators. We also claim that the generalized Hamiltonian Monte Carlo is preferable for achieving high sampling performance. The efficiency of the new methodology is assessed in comparison with state-of-the-art samplers, e.g. the No-U-Turn-Sampler, in real-world applications, such as endocrine therapy resistance in cancer, modeling of cell-cell adhesion dynamics and influenza A epidemic outbreak.
@@ 4.1 Randomization interval
GHMC s-AIA3 $\mathcal{U}\{1,...,1332\}$ $\mathcal{U}(0.036,0.052)$ $\mathcal{U}(0.0,0.5)$
@@
German HMC s-AIA3 $\mathcal{U}\{1,...,16\}$ $\mathcal{U}(0.111,0.160)$ 1.0
@@
GHMC s-AIA3 $\mathcal{U}\{1,...,7\}$ $\mathcal{U}(0.111,0.160)$ $\mathcal{U}(0.0,0.5)$
@@
Musk HMC s-AIA3 $\mathcal{U}\{1,...,110\}$ $\mathcal{U}(0.083,0.110)$ 1.0
@@
GHMC s-AIA3 $\mathcal{U}\{1,...,55\}$ $\mathcal{U}(0.084,0.121)$ $\mathcal{U}(0.0,0.5)$
@@ 4.2 Validation
The remaining HMC parameters, namely, a number of integration steps per Monte Carlo iteration, $L$, and numbers of burn-in and production iterations, are kept the same as in the numerical experiments in sAIA_paper. In addition, we conduct GHMC simulations to evaluate the efficacy of the proposed randomization interval with this sampler in comparison with HMC. For the PMU in Eq. (9), we set $\varphi$ to be uniformly selected from the interval $(0.0,0.5)$ (a standard recommendation) and use twice shorter trajectories than for HMC. We remark that this choice of GHMC parameters is based on heuristic intuition. The optimal selection of these parameters are investigated in the following sections.
@@ 8.2 Benchmark tests
Banana $\mathcal{U}\{2,5,7\}$ $\mathcal{U}(0.21904,1.00000)$
@@
G1000 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 39373.1 2387.2 1491.1 2509.4 5021.0
meanESS 0.5 532.1 1071.4 2142.4 4285.3
multiESS 0.4 299.8 402.5 895.4 3069.2
1 (HMC) minESS 132650.2 46712.1 21620.1 11740.0 5554.2
meanESS 2.5 533.6 1054.3 2093.2 4198.1
multiESS 3.7 301.3 1001.0 2000.0 4001.0
G500 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 31792.2 1553.4 919.8 1090.2 2339.1
meanESS 0.3 260.0 520.7 1043.2 2093.1
multiESS 0.3 170.6 320.8 384.0 1422.3
1 (HMC) minESS 116471.2 42111.1 21821.2 10094.4 4790.2
meanESS 2.4 261.3 515.5 1035.3 2046.0
multiESS 3.7 107.8 321.6 563.3 1325.2
G2000 $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 63451.2 3305.0 2515.3 5036.4 9195.1
meanESS 0.7 1104.1 2207.4 4417.2 8853.4
multiESS 0.5 428.5 860.1 1723.0 7096.3
1 (HMC) minESS 174435.1 47859.0 22770.3 10870.1 9582.1
meanESS 2.8 1099.1 2178.4 4348.1 8671.2
multiESS 3.7 677.3 2000.0 4001.0 8000.0
German $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 1.7 13.9 25.9 53.4 113.6
meanESS 0.3 10.2 23.8 49.1 101.0
multiESS 0.1 10.6 25.1 49.5 100.5
1 (HMC) minESS 24.9 13.3 26.2 53.1 112.6
meanESS 5.9 10.6 24.5 49.2 102.9
multiESS 2.3 10.9 25.4 49.4 100.8
Musk $\mathcal{U}(\varphi_{\text{lower}},\varphi_{\text{upper}})$ minESS 216.8 143.7 229.4 389.7 601.3 1458.3
meanESS 145.4 52.1 68.4 165.0 464.5 1031.2
multiESS 19.8 12.8 36.0 160.4 342.0 774.5
1 (HMC) minESS 6026.1 1156.4 187.7 290.3 525.7 1215.2
meanESS 3068.0 705.6 143.5 176.3 454.6 1047.4
multiESS 375.8 73.0 50.6 159.5 279.5 506.9
@@
AT-GHMC $\mathcal{U}(0.152,0.219)$ $\mathcal{U}(2,8)$ $\mathcal{U}(0.21904,1.00000)$ 0.92
HMC tuned $0.111$ 77 1.00000 0.69
HMC $0.050$ 22 $1.00000$ 0.91
$0.500$ 0.02
@@ 9.1 Patient resistance to endocrine therapy in breast cancer
NUTS 100 $\overline{L}=15$ $\overline{\Delta t}=0.247$ -
@@
NUTS 100 $\overline{L}=12$ $\overline{\Delta t}=0.354$ -
@@
NUTS 100 $\overline{L}=8$ $\overline{\Delta t}=0.474$ -
@@ 9.3 Influenza A (H1N1) epidemics outbreak
As an illustrative example, we consider an outbreak of influenza A (H1N1) in 1978 at a British boarding school involving 763 students. The dataset, available in the R package outbreaks outbreaks20, records the daily number of students confined to bed over a 14-day period. Following Grinsztajn2020, we model the outbreak dynamics using a classical SIR (Susceptible-Infected-Recovered) model, coupled with a negative binomial observation model to account for overdispersion in the reported case counts. The model includes the transmission rate, $\beta$, and the recovery rate, $\gamma$, which govern the transitions from susceptible to infected and from infected to recovered, respectively. For the prior distributions, we assume $\beta\sim\mathcal{N}^{0}(1,1)$ and $\gamma\sim\mathcal{N}^{0}(0.4,0.5)$, where $\mathcal{N}^{a}$ denotes a normal distribution truncated at $a$. The prior specification is completed by assigning an exponential prior to the inverse of the dispersion parameter, $\phi$, of the negative binomial distribution, i.e. $1/\phi\sim\text{Exp}(5)$. The initial conditions are set as $I_{0}=1$, $R_{0}=0$, and thus $S_{0}=762$.
@@
NUTS 140 $\overline{L}=5$ $\overline{\Delta t}=0.607$ -
@@ Appendix D Sensitivity analysis: Optimal $h_{\text{lower}}$
G1000 HMC (15) 5554.2 4198.1 4001.0
(15)  $5\%$ 5034.1 4361.2 4001.0
(15) + $5\%$ 5107.4 4483.2 4001.0
GHMC (15) 2518.3 2081.1 2000.0
(15)  $5\%$ 2818.4 2189.3 2000.0
(15) + $5\%$ 2772.3 2231.2 2000.0
German HMC (15) 26.2 24.5 25.4
(15)  $5\%$ 25.3 24.8 25.2
(15) + $5\%$ 25.4 25.0 25.3
GHMC (15) 12.4 10.1 10.7
(15)  $5\%$ 11.4 10.2 10.6
(15) + $5\%$ 11.3 10.2 10.7